# Extension principles for the Einstein Yang–Mills system

### v1

@@ 3.1.5. Pointwise estimates
$\mu\lesssim r^{b}\quad\text{and}\quad\lim_{u\rightarrow\underline{u}_{\ast}}\sup_{\mathcal{D}(u,\,\underline{u}_{\ast})}\frac{\mu}{r^{b}}=0.$
In particular, since $b\geq 0$, there is
$\mu|_{\Gamma}=0.$
@@
Since $0\leq r^{\prime}\leq r$, and hence $\frac{1}{r}\leq\frac{1}{r^{\prime}}$, the energy bound (3.1.4) leads to
$\frac{m}{r^{1+b}}\leq\int_{0}^{r}\frac{\left(\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)}{r^{\prime 1+b}}\mathrm{d}r^{\prime}|_{C_{u}}\leq E_{0}\lesssim I_{0}.$
@@
As a sequel, we will improve the pointwise estimates for $\alpha$ and $r^{-1}\hat{w}$.
@@
$\alpha^{2}\lesssim{}$ $r^{b}\quad\text{and}\quad\alpha^{2}\lesssim\varepsilon,$
$r^{-2}\hat{w}^{2}\lesssim{}$ $r^{b}\quad\text{and}\quad r^{-2}\hat{w}^{2}\lesssim\varepsilon^{2}.$
Remark 3.6.
If $b>0$, Lemma 3.5 suggests
$\alpha|_{\Gamma}=0\quad\text{and}\quad r^{-1}\hat{w}|_{\Gamma}=0.$
@@
[…]
where the bootstrap assumption (3.10) and the energy bound (3.1.4) are used. That is, we have derived a preliminary estimate for $\alpha$,
@@
[…]
Here in the last inequality, we have access to the equivalence (3.22) and the energy bound (3.1.4). As a result,
$\alpha^{2}\lesssim\mu+\mu^{\frac{1}{2}}r^{\frac{b}{2}}.$
Combined with the smallness of $\mu$ and Lemma 3.4, the first part of Lemma 3.5 is justified.
@@
[…]
@@
which entails $r\cdot\left(r^{-2}\hat{w}^{2}\right)|_{\Gamma}=0$. It enables us to integrate $\partial_{r}\left(r\cdot\left(r^{-2}\hat{w}^{2}\right)\right)$ in $r$ on each $C_{u}$ and then an application of the Cauchy-Schwarz inequality leads to
@@
[…]
@@
With Lemma 3.5 concluded, we have improved the pointwise estimates for $r^{-1}\hat{w}$ and $\alpha$ and hence close the bootstrap argument. In the subsequent analysis, we will derive more precise estimates to provide a detailed description for the solution.
3.1.6. Estimates for the geometry
@@
Before going into the pointwise estimates, we will deduce some integral bounds for $\mu$.
@@
[…]
@@
[…]
@@
and as a consequence of the energy bound (3.8),
@@
In view of Lemma 3.4, $01+h\underline{h}|_{\Gamma}=\mu|_{\Gamma}=$ 0, and hence $h|_{\Gamma}=-\underline{h}^{-1}|_{\Gamma}=(\Omega^{2}h)^{-1}|_{\Gamma}$, the estimate for $\ln h$ is proved.
In particular, if $b>0$, we can further show the solution $w\in C^{1}$.
Lemma 3.8.
In the case of $0<b<\frac{\sqrt{89}-9}{4}$, we have the following estimate for $\partial_{u}w$,
$|\partial_{u}w|^{2}\lesssim r^{b}.$
Proof.
The fact $u=\underline{u}$ on $\Gamma$ and the boundary condition $w|_{\Gamma}=\pm 1$, plus Remark 3.6 entail
$\partial_{u}w|_{\Gamma}=-\partial_{\underline{u}}w|_{\Gamma}=0.$
Now we can integrate the YM equation (2.5) from the axis and use the $L^{\infty}$ estimate for $\Phi$ (Lemma 3.5) to conclude
$|\partial_{u}w|\lesssim{}$ $|\partial_{u}w|_{\Gamma}+\int_{0}^{r}h^{-1}|w|\frac{|\Phi|}{r^{\prime}}\mathrm{d}r^{\prime}|_{C_{u}}$
[…]
3.1.7. Proof of Theorem 3.1
For the SSEYM system, the first singularity must emanate from the axis. One may refer to Theorem B.2 for more details. Therefore, it suffices to focus on a small neighbourhood of the tip $u=\underline{u}=\underline{u}_{\ast}$, and we may assume $u_{0}$ is close to $\underline{u}_{\ast}$ and
@@
As suggested by Lemma 3.2, the solution ($\alpha$) extends as an $H^{1}$ solution on the closure of the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$. To complete the proof Theorem 3.1, we have to show that the solution can be extended beyond $\underline{C}_{\underline{u}_{\ast}}$. More explicitly, we shall verify that the solution can be extended to $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$ if $\zeta>0$ is small enough.
@@
$\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\,\,\text{on}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}],$ (3.27)
@@
$u_{2}=\sup\,\{u_{1}\in(u_{0},\,\underline{u}_{\ast}]:\,\,\sup_{\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta)}\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\}.$
@@
$\mu<\frac{\varepsilon_{1}+\varepsilon}{2},\quad\text{on}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}]$
@@
Preliminarily, by virtue of the local existence theorem (Appendix A) and Lemma 3.2, the bootstrap assumption (3.27) leads to an $H^{1}$ solution on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$. In particular, the $H^{1}$ norm
@@
On any $C_{u}$, $u\in[u_{0},\,\underline{u}_{\ast}]$, an integration in $\underline{u}^{\prime}\in[\underline{u}_{\ast},\,\underline{u}]$ as (3.1.6) yields the identity
$\mu\Big{|}_{\underline{u}_{\ast}}^{\underline{u}}+\int_{r(u,\,\underline{u}_{\ast})}^{r(u,\,\underline{u})}\frac{\mu}{r}\mathrm{d}r|_{C_{u}}=\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},$
@@
$\mu(u,\,\underline{u})<\mu(u,\,\underline{u}_{\ast})+\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},\quad\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
@@
$\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime}<\frac{\varepsilon_{1}-\varepsilon}{2},\quad\forall\,\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
@@
In summary, we have proved (3.27) holds with $u_{1}=\underline{u}_{\ast}$. As a result, by the proof leading to Lemma 3.2, we know that the solution extends as an $H^{1}$ solution on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$. Combined with the local existence theorem (Appendix A) for $H^{1}$ data, it further allows us to extend the solution to $\mathcal{D}(\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta^{\prime})$ for some $\zeta^{\prime}\in(0,\,\zeta]$. We complete the proof of Theorem 3.1.
@@ 3.2. Cauchy stability in $H^{1}$ space
Having established the extension theorem 3.1, we will justify the Cauchy stability within the same topology. This amounts to proving a small data global existence in the finite domain $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$, where $\underline{u}_{\ast}\in(u_{0},\,+\infty)$. Recalling that the energies $E_{i}$, $i=0,1$, and $\underline{E}_{0}$, $\underline{E}_{0}^{\prime}$, are defined as in (3.1), we now state the Cauchy stability theorem.
@@
with $\varepsilon$ being small enough.
Case I: $0<b<\frac{\sqrt{89}-9}{4}$. There is a unique solution $\left(r,m,\alpha\right)\in C^{1}\times C^{1}\times H^{1}$ on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$ such that
$r^{-2}|\hat{w}|^{2}+|\alpha|^{2}+|\partial_{u}w|^{2}+\mu\lesssim\varepsilon^{2}r^{b},$
and for $u_{0}<u\leq\underline{u}_{\ast}$, $u_{0}<\underline{u}\leq\underline{u}_{\ast}$,
@@
Case II: $b=0$. There is a unique solution $\left(r,m,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$ on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$ such that
$r^{-2}|\hat{w}|^{2}+|\alpha|^{2}+\mu\lesssim\varepsilon^{2},$
and (3.29) holds as well.
@@
The proof goes through in line with the one leading to Theorem 3.1 and thus we omit the details.
4. The weighted $H^{1}$ case
Given two constants $u_{0}$ and $\underline{u}_{\ast}$ with $u_{0}<\underline{u}_{\ast}$, for any $\delta>0$, we will consider in the domain $\mathcal{D}(u_{0},\,u;\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$, a class of weighted energies. The weighted norm is defined through a scale invariant function $\xi$ (1.8). Indeed, we will prove later in Proposition 4.4 that in the domain concerned,
$10\leq\xi<$ 1 .
Before the statement of the main theorem, let us introduce some additional notations for this section. We denote the strip
$\mathcal{S}(u,\,\underline{u}):=[u_{0},\,u]\times[\underline{u}_{\ast},\,\underline{u}],\quad u\in(u_{0},\,\underline{u}_{\ast}),\,\,\underline{u}\in(\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta],$ (4.1)
and $\mathcal{S}(u,\,\underline{u})$ will be simplified as $\mathcal{S}$. Besides, we abbreviate $C_{u}^{[\underline{u}_{\ast},\,\underline{u}]}$ by $C_{u}$, and $\underline{C}_{\underline{u}}^{[u_{0},\,u]}$ by $\underline{C}_{\underline{u}}$. We also denote the value of $m$ at $\underline{u}=\underline{u}_{\ast}$ by $m_{\ast}$. Similar conventions apply to $\underline{h}_{\ast},\,\mu_{\ast},\cdots$.
Fix $b$ as in (1.4), and a constant $\gamma\in(b,2]$. We define the $\xi$ weighted energy norms
$\|\alpha\|^{2}_{H^{0}_{w}(C_{u})}:={}$ $\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(\frac{r^{-2}\hat{w}^{2}}{\xi^{3+\gamma}}+\frac{\alpha^{2}}{\xi^{1+\gamma}}\right)\mathrm{d}\underline{u},$ (4.2a)
$\|\hat{w}\|^{2}_{H^{1}_{w}(\underline{C}_{\underline{u}})}:={}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\frac{(\partial_{u}w)^{2}}{\xi^{3+\gamma}}+\frac{r^{-2}\hat{w}^{2}}{\xi^{1+\gamma}}\right)\mathrm{d}u,$ (4.2b)
$\|\alpha\|^{2}_{H^{1}_{w}(C_{u})}:={}$ $\|\alpha\|^{2}_{H^{0}_{w}(C_{u})}+\int_{C_{u}}\xi^{1-\gamma}\frac{(\alpha^{\prime})^{2}}{r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u},$ (4.2c)
$\|\hat{w}\|^{2}_{H^{1^{\prime}}_{w}(\underline{C}_{\underline{u}})}:={}$ $\|\hat{w}\|^{2}_{H^{1}_{w}(\underline{C}_{\underline{u}})}+\int_{\underline{C}_{\underline{u}}}\xi^{1-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u.$ (4.2d)
4.1. Extension theorem in the weighted $H^{1}$ space

### v2

@@ 3.6. Pointwise estimates
$\lim_{u\rightarrow\underline{u}_{\ast}}\sup_{\mathcal{D}(u,\,\underline{u}_{\ast})}\mu=0,\quad\text{and}\quad\mu|_{\Gamma}=0.$
@@
Since $0\leq r^{\prime}\leq r$, and hence $\frac{1}{r}\leq\frac{1}{r^{\prime}}$, the energy bound (3.5) leads to
$\frac{m}{r}\leq\int_{0}^{r}\frac{\left(\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)}{r^{\prime}}\mathrm{d}r^{\prime}|_{C_{u}}\leq E_{0}\lesssim I_{0}.$
@@
As a next step, we will improve the pointwise estimates for $\alpha$ and $r^{-1}\hat{w}$.
@@
$r^{-2}\hat{w}^{2}\lesssim\varepsilon^{2}\quad\text{and}\quad\alpha^{2}\lesssim\varepsilon,$
and
$r^{-2}\hat{w}^{2}|_{\Gamma}=0\quad\text{and}\quad\alpha^{2}|_{\Gamma}=0.$
@@
$\lesssim{}$ $\underline{E}_{0}^{\prime}\lesssim I_{1},$
where the bootstrap assumption (3.8) and the energy bound (3.5) are used. That is, we have derived a preliminary estimate for $\alpha$,
@@
$\leq{}$ $\int_{0}^{r}\alpha^{2}\mathrm{d}r^{\prime}|_{C_{u}}+\left(\int_{0}^{r}\alpha^{2}\mathrm{d}r^{\prime}|_{C_{u}}\right)^{\frac{1}{2}}\left(\int_{0}^{r}(\alpha^{\prime})^{2}\mathrm{d}r^{\prime}|_{C_{u}}\right)^{\frac{1}{2}}$
$\lesssim{}$ $m+m^{\frac{1}{2}}r^{\frac{1}{2}}I_{1}^{\frac{1}{2}}.$
Here in the last inequality, we have access to the equivalence (3.20) and the energy bound (3.5). As a result,
$\alpha^{2}\lesssim\mu+\mu^{\frac{1}{2}}.$
Combined with Lemma 3.4, the first part of Lemma 3.5 is justified.
@@
$\lesssim{}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(h^{2}(\partial_{u}w)^{2}+\frac{\hat{w}^{2}}{r^{2}}\right)\mathrm{d}u\lesssim\underline{E}_{0}\lesssim I_{0}.$
@@
which further entails $r\cdot\left(r^{-2}\hat{w}^{2}\right)|_{\Gamma}=0$. It enables us to integrate $\partial_{r}\left(r\cdot\left(r^{-2}\hat{w}^{2}\right)\right)$ in $r$ on each $C_{u}$ and then an application of the Cauchy-Schwarz inequality leads to
@@
$\lesssim{}$ $\int_{0}^{r}\left(2\alpha^{2}+r^{-2}\hat{w}^{2}\right)\mathrm{d}r^{\prime}|_{C_{u}}\lesssim m.$
@@
With Lemma 3.5 concluded, we have improved the pointwise estimates for $r^{-1}\hat{w}$ and $\alpha$, thereby closing the bootstrap argument. In the subsequent analysis, we will derive more precise estimates to provide a detailed description for the solution.
3.7. Estimates for the geometry
@@
Before going into the pointwise estimates, we deduce some integral bounds for $\mu$.
@@
$={}$ $\int_{\underline{C}_{\underline{u}}}\left(2h^{2}\frac{(\partial_{u}w)^{2}}{r}+\left(1-\mu\right)\frac{\Phi^{2}}{r}\right)h^{-1}\mathrm{d}u\lesssim I_{0},$ (3.21)
@@
$={}$ $\int_{0}^{r}\left(2\left(1-\mu\right)\frac{\alpha^{2}}{r^{\prime}}+\frac{\Phi^{2}}{r^{\prime}}\right)\mathrm{d}r^{\prime}|_{C_{u}}\lesssim I_{0},$ (3.23)
@@
and as a consequence of the energy bound, there is (3.6),
@@
In view of Lemma 3.4, $1+h\underline{h}|_{\Gamma}=\mu|_{\Gamma}=0$, and hence $h|_{\Gamma}=-\underline{h}^{-1}|_{\Gamma}=(\Omega^{2}h)^{-1}|_{\Gamma}$, the estimate for $\ln h$ is proved. ∎
3.8. Proof of Theorem 3.1
For the SSEYM system, the first singularity must emanate from the axis. For further details, we refer to Theorem A.2. It therefore suffices to focus on a small neighbourhood of the tip $u=\underline{u}=\underline{u}_{\ast}$. We may assume $u_{0}$ is close to $\underline{u}_{\ast}$ and that
@@
As indicated by Lemma 3.2, the solution ($\alpha$) extends as an $H^{1}$ function to the closure of domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$. To complete the proof Theorem 3.1, it remains to show that the solution can be extended beyond $\underline{C}_{\underline{u}_{\ast}}$. Specifically, we must demonstrate that the solution can be extended to $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$ for some sufficiently small $\zeta>0$.
@@
$\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\,\,\text{in}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}],$ (3.25)
@@
$u_{2}=\sup\,\left\{u_{1}\in(u_{0},\,\underline{u}_{\ast}]:\,\,\sup_{\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta)}\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\right\}.$
@@
$\mu<\frac{\varepsilon_{1}+\varepsilon}{2},\quad\text{in}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}],$
@@
Preliminarily, by virtue of the local existence theorem (Section 4) and Lemma 3.2, the bootstrap assumption (3.25) leads to an $H^{1}$ solution in $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$. In particular, the $H^{1}$ norm
@@
On any $C_{u}$, $u\in[u_{0},\,\underline{u}_{\ast}]$, an integration in $\underline{u}^{\prime}\in[\underline{u}_{\ast},\,\underline{u}]$ as (3.7) yields the identity
$\mu\Big{|}_{\underline{u}_{\ast}}^{\underline{u}}+\int_{r(u,\,\underline{u}_{\ast})}^{r(u,\,\underline{u})}\frac{\mu}{r}\mathrm{d}r|_{C_{u}}=\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(2\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},$
@@
$\mu(u,\,\underline{u})<\mu(u,\,\underline{u}_{\ast})+\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(2\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},\quad\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
@@
$\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime}<\frac{\varepsilon_{1}-\varepsilon}{2},\quad\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
@@
In summary, we have proved (3.25) holds with $u_{1}=\underline{u}_{\ast}$. Consequently, following the argument leading to Lemma 3.2, the solution extends as an $H^{1}$ solution to the domian $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$. Together with the local existence theorem (Section 4) for $H^{1}$ data, it allows us to further extend the solution to $\mathcal{D}(\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta^{\prime})$ for some $\zeta^{\prime}\in(0,\,\zeta]$. This completes the proof of Theorem 3.1.
@@ 3.9. Cauchy stability in $H^{1}$ space
Having established the extension theorem (Theorem 3.1), we now turn to Cauchy stability within the same topology. This amounts to proving a small-data global existence on the finite domain $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$, where $\underline{u}_{\ast}\in(u_{0},\,+\infty)$. Recall that the energies $E_{i}$, $i=0,1$, and $\underline{E}_{0}$, $\underline{E}_{0}^{\prime}$, are defined as in (3.1). We now state the Cauchy stability theorem.
@@
with $\varepsilon$ sufficiently small. Then there exists a unique and global solution $\left(r,m,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$ in the domian $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$ such that
$r^{-2}|\hat{w}|^{2}+|\alpha|^{2}+\mu\lesssim\varepsilon^{2},$
and for $u_{0}<u\leq\underline{u}_{\ast}$, $u_{0}<\underline{u}\leq\underline{u}_{\ast}$, the solution satisfies the following unifom energy bound
@@
The proof goes through in line with the one leading to Theorem 3.1, and we therefore omit the details here.
4. Local existence theorem
For the characteristic initial value problem with data prescribed on a future light cone emanating from the center of symmetry, the local existence theorem for classical solutions to the spherically symmetric Einstein scalar field system was proved in [8]. Regarding to the proof, Christodoulou reduced the whole system in Bondi coordinates to a scalar equation and utilized the contracting mapping principle to obtain the existence theorem.
In this section, however, we aim to present a proof based on energy methods. In more details, we work in the double null gauge and demonstrate a local existence theory for the SSEYM system in $H^{1}$, using an iteration method.
The Yang–Mills equation can be alternatively written in terms of $\hat{w}$ (1.3)
$\partial_{u}\partial_{\underline{u}}\hat{w}+2\frac{\Omega^{2}}{r^{2}}\hat{w}=\Omega^{2}\frac{\hat{w}^{2}}{r^{2}}\left(\hat{w}+3\right).$ (4.1)